### v1

# Enhancing Outlier Knowledge for Few-Shot Out-of-Distribution Detection with Extensible Local Prompts

Abstract
Out-of-Distribution (OOD) detection, aiming to distinguish outliers from known categories, has gained prominence in practical scenarios. Recently, the advent of vision-language models (VLM) has heightened interest in enhancing OOD detection for VLM through few-shot tuning. However, existing methods mainly focus on optimizing global prompts, ignoring refined utilization of local information with regard to outliers. Motivated by this, we freeze global prompts and introduce a novel coarse-to-fine tuning paradigm to emphasize regional enhancement with local prompts. Our method comprises two integral components: global prompt guided negative augmentation and local prompt enhanced regional regularization. The former utilizes frozen, coarse global prompts as guiding cues to incorporate negative augmentation, thereby leveraging local outlier knowledge. The latter employs trainable local prompts and a regional regularization to capture local information effectively, aiding in outlier identification. We also propose regional-related metric to empower the enrichment of OOD detection. Moreover, since our approach explores enhancing local prompts only, it can be seamlessly integrated with trained global prompts during inference to boost the performance. Comprehensive experiments demonstrate the effectiveness and potential of our method. Notably, our method reduces average FPR95 by 5.17% against state-of-the-art method in 4-shot tuning on challenging ImageNet-1k dataset, even outperforming 16-shot results of previous methods.
1 Introduction
Out-of-distribution (OOD) detection [1, 2, 3, 4] aims to distinguish outliers, i.e., samples that do not belong to known in-distribution (ID) classes. It is crucial for industries that require a high level of safety [5], such as face recognition [6] and autonomous driving [7]. Most previous methods in the field of OOD detection handle the problem using single-modal methods and concentrate on post-hoc processing [8, 9] or leveraging outliers [10, 11]. However, they suffer from computational inefficiency or consuming data collection. Recently, as the emergence of vision-language models [12, 13] shows promising results in multi-modal tasks including image caption [14], video understanding [15], and so on, it is promising to exploit textual representations to improve performance of OOD detection given a vision-language model [12] as a prior model. Considering that zero-shot undergoes a domain gap between upstream and target distribution while full-tuning may pose threats to the learned representation, it is meaningful to explore few-shot learning for vision-language model in OOD detection, in which the detector is prohibited from real outliers and only has access to several ID images.
The most challenging scene for OOD detection is that one hard OOD sample is similar to a known class on the whole and only has subtle differences locally, which naturally requires the detector to identify outliers through local outlier regions. However, existing research falls short of refining OOD task via rich local information when subtle OOD samples are exhibited in certain regions, as is shown in Fig. 1. Some methods merely focus on utilizing global features only [16], which ignores local features and inevitably brings about coarse description. Others use the same prompts to match both global and local image features [17, 18], so the gap between them may lead to inaccurate local outlier identification. Consequently, it is straightforward that enhancing regional information to empower the model with local outlier knowledge could be significant to OOD detection.
Figure 1: Comparison of prompt learning for OOD detection task. Prompts with global optimization may fail in challenging OOD samples as they are overall similar to ID samples and only have subtle regional differences. For example, cat and tiger are generally similar (blue boxes) and only differ in forehead (red box). Our approach with local outlier knowledge tackles the issue to some extent.
Motivated by the above observations, we focus on how to explicitly exploit regional-related knowledge, i.e., the learned knowledge is determined by the position of regions, about not only known classes but also unknown samples, which has not been explored before. To tackle the issue, we focus on enhancing local features solely and propose the extensible local prompts for fine local information utilization. Concretely, we decompose global and local prompts for respective features, and propose two integral components: global prompt guided negative augmentation and local prompt enhanced regional regularization. Global prompt is frozen to provide guidance for generating negative augmentation, which could facilitate the exploitation of local outlier knowledge. Fine local prompt is designed for regional regularization to refine regional-related local prompts, which captures the fine local information to better distinguish OOD inputs. Moreover, we also propose corresponding regional-related evaluation metrics to leverage the power of local prompts.
Another advantage of our approach is that it orthogonally explores the benefits of local prompts shown in Fig. 1 and uses hand-crafted global prompts, i.e., a photo of a {class}. Hence, it is extensible to numerous global prompt designing methods [18, 19]. Therefore, we can extend it by seamlessly replacing coarse global prompts with trained ones during inference to improve the performance.
Comprehensive experiments demonstrate the effectiveness of the proposed method. Specifically, our results outperform the state-of-the-art methods in both OOD detection (5.17% reduction on FPR95) and ID accuracy evaluation (1.02%) under 4-shot tuning by a large margin, exhibiting the superiority and the potential of enhancing local prompts for OOD detection and ID classification ability. We summarize our contributions as follows:
  • We focus on enhancing local information with local prompts and propose a coarse-to-fine few-shot tuning paradigm. Specifically, we propose effective negative augmentation and regional regularization to learn local prompts with local outlier knowledge.
  • We propose regional-related OOD score and ID classification metrics, which benefit from enhanced local prompts in OOD detection. Moreover, our method is orthogonal to global prompt optimization methods and is extensible to get notable performance when integrated with trained local global prompts during inference.
  • We conduct comprehensive experiments and achieve competitive results on various OOD detection datasets, demonstrating the superiority of our method. Notably, our 4-shot results on challenging ImageNet-1k even outperform 16-shot results of previous methods.
2 Related work
2.1 Out-of-distribution detection with vision-language model
Out-of-distribution detection has been widely studied in the past decades [10, 20, 21, 11]. With the emergence of CLIP [12], which learns outstanding visual and textual representations pre-trained on large-scale image-text pairs and displays excellent ability in various downstream tasks, much attention has been paid to transferring vision-language model to OOD detection task [22, 16, 23, 19, 18, 24]. For instance, MCM [16] adopts the visual and textual representations extracted from CLIP encoders and applies softmax with temperature to better separate ID and OOD data. ZOC [23] generates pseudo unseen class labels with additional modules and defines a novel confidence score accordingly. Clipn [19] enhances the discrimination of OOD samples with extra encoder and data [25].
2.2 Multi-modal few-shot prompt learning
Since vision-language model like CLIP [12] achieves superior performance in numerous image-text tasks, many attempts have been made to explore few-shot tuning to leverage the power of large pre-trained models. There are mainly two lines of approaches, namely Adapter Learning [26, 27, 28, 29, 30] and Prompt Learning [31, 32]. Both of them froze the pre-trained encoders when fine-tuning a few additional parameters.
Prompt learning promotes the performance of vision-language model by enhancing the prompts in both visual [33, 32] and textual [31, 34] aspects. A common practice is setting textual prompts learnable and forwarding them together with visual inputs. CoOp [31] replaces the hand-crafted prompts with contextual learnable embeddings and optimizes them during few-shot training. In our approach, we freeze global prompts and fine-tune local prompts individually.
2.3 Global and local information
In Vision Transformer [35], global tokens are specifically designed for image classification task [36, 37]. They represent overall characteristics of images. Local tokens are utilized for dense prediction tasks including segmentation [38, 39], object detection [40, 41], and so on. The utilization of local features has been explored as well, such as pyramid features [42] and window attention [43, 44, 45].
In the field of OOD detection, global features are primarily employed and numerous studies have been conducted concerning local information [17, 18, 46]. GL-MCM [17] takes all tokens into consideration and uses the sum of global and local OOD scores to measure the confidence. LoCoOp [18] keeps global ID textual embeddings away from the interference of ID-irrelevant regions. However, all existing approaches take same prompts for all features. By contrast, our method directly enhances OOD detection with ID-related areas and refines local prompts to leverage local outlier knowledge.
3 Preliminary
3.1 Problem definition
Formally, out-of-distribution detection can be viewed as a binary classification that the detector has to identify whether the input image is from ID or OOD space. Labeled training data is composed of $\mathcal{D}^{\mathrm{in}}_{\mathrm{train}}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, where $\bm{x}_{i}$ is sampled i.i.d. from joint data distribution space $\mathcal{P}_{\mathcal{XY}^{\mathrm{in}}}$ and $y_{i}\in\mathcal{Y}^{\mathrm{in}}$. $\mathcal{Y}^{\mathrm{in}}=\{1,\cdots,C\}$ is the label space of ID data and $C$ is the number of classes. We transfer the definition of few-shot learning in classification to OOD detection to fine-tune given vision-language model instead of training from scratch. Under few-shot setting, a small proportion of images in each class (e.g., $1$, $4$, or $16$ images) are extracted for training. $\mathcal{Y}^{\mathrm{out}}$ is the label space of OOD data and there is no overlap between ID and OOD label, i.e., $\mathcal{Y}^{\mathrm{in}}\cap\mathcal{Y}^{\mathrm{out}}=\emptyset$. Typically, the model is unable to obtain any OOD data during training process.
The test dataset is a mixture of $\mathcal{D}^{\mathrm{in}}_{\mathrm{test}}$ and $\mathcal{D}^{\mathrm{out}}_{\mathrm{test}}$. When testing, OOD performance is evaluated by discriminating which distribution each sample comes from. A discriminant function is calculated to identify ID and OOD samples:
$D(\bm{x})=\left\{\begin{aligned} \mathrm{ID},&\qquad S(\bm{x})\geq\gamma\\ \mathrm{OOD},&\qquad S(\bm{x})<\gamma\\ \end{aligned}\right.,$ (1)
where $S(\bm{x})$ is the confidence score function measuring the uncertainty and $\gamma$ is the threshold.
3.2 OOD detetion with CLIP
Feature extraction
CLIP [12] is compromised of image encoder $E_{\mathrm{I}}$ and text encoder $E_{\mathrm{T}}$. Given an image $I$ and a text $T$ describing the corresponding image, they are extracted into feature embeddings, respectively. Specifically, text can be hand-crafted templates [12] or learnable context words [34], i.e., $\bm{t}_{c}=[\bm{v}_{1},\cdots,\bm{v}_{L};\bm{v}_{c}]$, where $\bm{v}_{i},(i=1,\cdots,L)$ is the learnable embedding, and $L$ is the length of context words. $\bm{v}_{c}$ is the embedding of class name. Text feature is obtained from text encoder $\bm{t}=E_{\mathrm{T}}(\bm{t}_{c}):\mathbb{R}^{(L+1)\times d}\rightarrow\mathbb{R}^{d}$.
Image is first split into several patches with an additional class token embedding $\bm{x}\in\mathbb{R}^{(N+1)\times d}$. $N$ and $d$ denote the number of local tokens and hidden dimension, respectively. Both global and local visual features are then extracted from image encoder $[\bm{z}^{\mathrm{g}},\bm{z}^{\mathrm{l}}]=E_{\mathrm{I}}(\bm{x}):\mathbb{R}^{(N+1)\times d}\rightarrow\mathbb{R}^{(N+1)\times d}$.
OOD score calculation
The MCM score [16] is defined as the maximum of similarity after $\mathrm{softmax}$ with temperature:
$S_{\mathrm{MCM}}(\bm{x})=\max_{i}\frac{\mathrm{exp}(\mathrm{sim}(\bm{z}^{\mathrm{g}},\bm{t}_{i})/T)}{\sum_{j=1}^{C}{\mathrm{exp}(\mathrm{sim}(\bm{z}^{\mathrm{g}},\bm{t}_{j})/T)}},$ (2)
and GL-MCM score [17] considers local information. It applies softmax for local features as well and the maximum is selected as local MCM score. Final OOD score is the sum of MCM score and local MCM score:
$S_{\mathrm{GL-MCM}}(\bm{x})=S_{\mathrm{MCM}}(\bm{x})+\max_{i,h}\frac{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)}{\sum_{j=1}^{C}{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{j})/T)}},$ (3)
where $\bm{z}^{\mathrm{g}}$ represents global feature and $\bm{z}^{\mathrm{l}}_{h}~{}(h\in\{1,\cdots,N\})$ represents all extracted local features.
4 Methodology
We propose a coarse-to-fine paradigm to strengthen OOD detection with local outlier knowledge. We first present negative augmentation that leverages local and unknown outlier information. Then, we describe local prompt enhancement along with their training regularization and a regional OOD score for better regional information utilization. Finally, we extend our method to trained global prompts to boost the performance. Detailed structures are shown in Fig. 2.
4.1 Global Prompt Guided Negative Augmentation
As it is costly and not always effective to rely solely on real outliers, one crucial problem for OOD detection is how to leverage imaginary local outlier knowledge to the detector. We propose to synthesize hard negative images to force the model to learn features with strong relation only to simulate outlier situations while avoiding the requirement of real OOD samples. To this end, we apply a simple and straightforward random crop augmentation to empower the detector with unknown (i.e., OOD) information. Specifically, we randomly crop image input $m$ times and use the hand-crafted template with corresponding class name, i.e., $\bm{t}_{c}=$ “a photo of {class}"""" as text inputs to calculate the image-text similarity with global features from image encoder. We then select $m_{1}$/$m_{2}$ images with the largest/least similarity, respectively. Images with the largest similarity can be seen as positive samples and those with the least similarity serve as hard negative samples. Notably, random crop is helpful to leverage local information in that local prompts can learn regional-related outlier information through space translation by means of random crop. Effectiveness of random crop is verified in Sec. 5.3.
Figure 2: Detailed structure of the proposed regional enhancement framework. Our method consists of global prompt guided negative augmentation and local prompt enhanced regional regularization. We froze global prompts to select regional augmented samples and enhance local prompts to learn regional-related representation that helps improve both ID accuracy and OOD detection.
Global prompts can be viewed as coarse guidance for negative augmentation standing for overall representation and are frozen in our framework. Motivated by previous observations[16, 17], we simply use the basic global prompts during training and select augmented images as they are fairly good for extracting global features. Concretely, the global prompts are used to (1) guide the negative augmentation selection, which has been described above; (2) guide the global OOD score calculation for evaluation, as shown in Sec. 4.3.
It is worth emphasizing that our approach is orthogonal to all existing global prompt optimization strategies, i.e., global prompts are built without tuning in our structure [18], so during inference, it is expected that extending the hand-crafted global prompts with carefully designed ones can promote the performance. Concretely, we replace hand-crafted global prompts with trained ones that have the same shape. Note that our main purpose is to decompose global and local prompts and showcase the effectiveness of local outlier enhancement for OOD detection. Therefore, we do not specifically select the template as it is not the focus of the paper.
4.2 Local prompt enhanced regional regularization
Once augmented inputs are decided, corresponding local features are then used to optimize local prompts to identify ID and OOD samples from the perspective of regional enhancement. We aim at utilizing the local features extracted from image encoder in a fine way to (1) learn regional-related prompts that better characterize local similarity for both ID and OOD regions; (2) enhance detection ability with refined local features. Specifically, we design local prompts with learnable context words $\bm{t}_{\mathrm{in}}$ to represent local textual information for each known class. Moreover, on account of the existence of OOD regions, we build a few local negative prompts $\bm{t}_{\mathrm{neg}}$ to handle possible local outliers.
Contrary to previous global prompt optimization methods, local prompts that work on local features are expected to learn regional-related textual representations and be aware of outliers during training. To this end, we enhance local prompts through regional regularization, which consists of contrastive regularization and diversity regularization, to better distinguish between ID and OOD samples.
Contrastive regularization
We propose regional contrastive regularization to enhance local textual prompts for better local information utilization. It is composed of local loss and local negative loss. Given a positive sample of training image, local loss is employed to learn regional information of known classes:
$\mathcal{L}_{\mathrm{pos}}(\bm{x},y)=-\mathrm{log}\frac{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)\}}{\sum_{\mathrm{in}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{in}})/T)\}}+\sum_{\mathrm{neg}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{neg}}})/T)\}},$ (4)
where $(\bm{x},y)$ stands for training image-label pairs, $\mathcal{T}_{k}(\bm{x})$ is the sum of $k$ largest elements in $\bm{x}$, $k$ is regional number and $T$ is the temperature. We use cosine similarity as the similarity metric:
$\mathrm{sim}(\bm{t}_{i},\bm{t}_{j})=\frac{\bm{t}_{i}\cdot\bm{t}_{j}}{\left\|{\bm{t}_{i}}\right\|\left\|{\bm{t}_{j}}\right\|}.$ (5)
Similarly, given a hard negative sample of the image denoted as $\tilde{\bm{x}}$, local negative loss is defined to make the model aware of outliers:
$\mathcal{L}_{\mathrm{neg}}(\tilde{\bm{x}})=-\mathrm{log}\frac{\sum_{\mathrm{neg}}\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{neg}})/T)\}}{\sum_{\mathrm{in}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{in}})/T)\}}+\sum_{\mathrm{neg}}{\mathcal{T}_{k}\{\mathrm{exp}(\mathrm{sim}(\bm{\tilde{z}}_{h}^{\mathrm{l}},\bm{t}_{\mathrm{neg}})/T)}\}}.$ (6)
Intuitively, contrastive regularization forces the model to focus on ID-related regions through random crop, and keep away from outlier-related regions through negative augmentation. We demonstrate the necessity of local negative loss to make margins for unknown categories in OOD detection in Sec.5.3.
Diversity regularization
As local negative prompts are randomly initialized, more regularization has to be imposed on them to ensure their diversity. Therefore, in addition to the proposed contrastive regularization, we apply a diversity regularization on local negative prompts, as follows:
$\mathcal{L}_{\mathrm{reg}}=\frac{1}{N_{\mathrm{neg}}^{2}}\sum_{1\leq i,j\leq N_{\mathrm{neg}}}\mathrm{sim}(\bm{t}_{i},\bm{t}_{j}),$ (7)
where $N_{\mathrm{neg}}$ is the number of local negative prompts. Final Loss is a weighted sum of the above loss:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{pos}}+\lambda_{\mathrm{neg}}\mathcal{L}_{\mathrm{neg}}+\lambda_{\mathrm{reg}}\mathcal{L}_{\mathrm{reg}},$ (8)
where $\lambda_{\mathrm{neg}}$ and $\lambda_{\mathrm{reg}}$ are the coefficients of the corresponding losses.
4.3 Regional OOD score
On consideration that regional information is especially enhanced during training, we propose Regional-MCM score to enhance OOD score beyond simple MCM and GL-MCM:
$S_{\mathrm{R-MCM}}(\bm{x})=S_{\mathrm{MCM}}(\bm{x})+\mathcal{T}_{k}^{\mathrm{mean}}\{\frac{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)}{\sum_{j=1}^{C+N_{\mathrm{neg}}}{\mathrm{exp}(\mathrm{sim}(\bm{z}_{h}^{l},\bm{t}_{j}})/T)}\},$ (9)
where $\mathcal{T}_{k}^{\mathrm{mean}}(\bm{x})$ is the mean of $k$ largest elements in $\bm{x}$ and $k$ represents regional number. Intuitively, Regional-MCM serves as a general form of GL-MCM that takes the $k$ most similar regions into consideration and additionally contains local negative prompts, which is helpful to consider more than one certain region with high similarity and improve OOD detection ability. Effectiveness of the score is verified in Sec. 5.3.
As the proposed OOD score is only applicable for OOD detection, we further propose a local-aware score that assigns a score to each class and thus determines ID category with the largest score. Specifically, ID classification takes both global and local features into consideration:
$h_{i}^{\mathrm{g}}(\bm{x})=\rm sim(\bm{z}^{\mathrm{g}},\bm{t}_{i}),$ (10)
$h_{i}^{\mathrm{l}}(\bm{x})=\mathcal{T}_{k}^{\mathrm{mean}}\{\rm exp(sim(\bm{z}_{h}^{\mathrm{l}},\bm{t}_{i})/T)\},$ (11)
where $h_{i}^{\mathrm{g}}(\bm{x})$, $h_{i}^{\mathrm{l}}(\bm{x})$ are global and local portions and $i$ represents the category label. The score attached to each category is measured by:
$f(y=i|\bm{x})=h_{i}^{\mathrm{g}}(\bm{x})*h_{i}^{\mathrm{l}}(\bm{x}),$ (12)
each sample is then classified corresponding to the maximum score.
[…]
Method iNaturalist SUN Places Texture Average
[…]
Zero-shot
MCM [16] 30.91 94.61 37.59 92.57 44.69 89.77 57.77 86.11 42.74 90.77
GL-MCM [17] 15.18 96.71 30.42 93.09 38.85 89.90 57.93 83.63 35.47 90.83
Full-tuning
MSP [8] 54.05 87.43 73.37 78.03 72.98 78.03 68.85 79.06 67.31 80.64
Energy [47] 29.75 94.68 53.18 87.33 56.40 85.60 51.35 88.00 47.67 88.90
KNN [48] 29.17 94.52 35.62 92.67 39.61 91.02 64.35 85.67 42.19 90.97
NPOS [4] 16.58 96.19 43.77 90.44 45.27 89.44 46.12 88.80 37.93 91.22
Textual-OE [24] 29.61 94.74 57.12 87.34 66.82 83.71 79.29 77.76 58.21 85.88
Few-shot    4-shot
CoOp [31] 18.95 95.52 29.58 92.90 38.72 89.64 48.03 85.87 33.82 90.98
Ours 9.65 97.87 20.40 95.57 29.39 92.67 51.20 88.00 27.66 93.53
LoCoOp [18] 21.67 95.69 22.98 95.07 31.41 92.10 49.79 87.85 31.46 92.68
Ours+LoCoOp 12.81 97.29 19.34 95.85 27.53 92.97 45.51 89.99 26.29 94.03
   16-shot
CoOp [31] 14.60 96.62 28.48 92.65 36.49 89.98 43.13 88.03 30.67 91.82
Ours 8.71 98.10 23.97 94.85 32.50 92.32 47.93 89.04 28.27 93.58
LoCoOp [18] 16.05 96.86 23.44 95.07 32.87 91.98 42.28 90.19 28.66 93.52
Ours+LoCoOp 8.63 98.07 23.23 95.12 31.74 92.42 34.50 92.29 24.52 94.48
Table 2: ID accuracy evaluation using ImageNet-1k as ID data.
Method Tuning Manner ID accuracy
MCM [16] zero-shot 67.01
GL-MCM [17] 67.01
KNN [48] full-tuning 79.64
NPOS [4] 79.42
LoCoOp [18] 4-shot 69.50
Ours 70.52
LoCoOp [18] 16-shot 71.86
Ours 74.24
Table 3: Comparison results of near OOD detection
tasks. We report 16-shot results for few-shot methods.
Our method can achieve performance against other methods.
Method ID ImageNet-10 ImageNet-20 Average
OOD ImageNet-20 ImageNet-10
[…]
MCM [16] 5.00 98.71 12.51 97.70 8.75 98.21
GL-MCM [17] 10.10 98.04 9.00 98.62 9.55 98.33
LoCoOp [18] 11.20 97.49 12.00 97.79 11.60 97.64
Ours 3.90 99.06 6.20 98.84 5.05 98.95
5 Experiments
5.1 Experimental setup
Datasets
Following existing works [16, 17, 18], large scale ImageNet-1K [49] along with a 100-category subset of it denoted as ImageNet-100 are used as ID dataset and OOD dataset is a combination of iNaturalist [50], SUN [51], Places [52] and Texture [53]. In addition, We also use two semantically similar subsets of ImageNet-1k, i.e., ImageNet-10 and ImageNet-20, to evaluate near OOD detection performance [16].
Implementation details
We use CLIP-Base/16 as the backbone. Image encoder, text encoder, and global prompts are frozen and only local prompts are learnable. We set one local prompt for each known category (1000 in total for ImageNet-1k). $\lambda_{\mathrm{neg}}$ and $\lambda_{\mathrm{reg}}$ are $5$ and $0.5$, respectively. As the stability of temperature $T$ has been verified by previous research [16, 17], we set $T$ to be 1 by default.
Evaluation metrics
We report the following metrics for evaluation: (1) the area under the receiver operating characteristic curve (AUROC); (2) false positive rate of OOD samples when true positive rate of ID samples is 95% (FPR95); (3) in-distribution data classification accuracy (ID accuracy).
5.2 Main results
ImageNet-1k as ID dataset
We report 4 and 16-shot results on four common OOD datasets using ImageNet-1k as ID dataset. It can be seen in Tab. 1 that tuning local prompts only using hand-crafted global prompts (Ours) can achieve competitive results, especially in datasets with challenging fine local textures like iNaturalist. Concretely, we get impressive progress in 4-shot tuning (outperforming by 3.80% on FPR95). In 16-shot setting, our approach gets competitive results as well. All the results above strongly showcase the potential of regional enhancement for OOD detection as an orthogonal direction to global prompt optimization methods.
Figure 3: Comparison of density map on iNaturalist. Ours are more separable.
Extensive to global prompts
Benefiting from the convenient integration with well-trained global prompts, we combine the advantage of trained global prompts from LoCoOp [18] without further design (Ours+LoCoOp) and once again get substantial improvements on all datasets. For example, we surpass the state-of-the-art method by a large margin (5.17% on FPR95 and 1.35% on AUROC), even against previous methods with 16 shots. Specifically, extending local prompts to global prompts is helpful to datasets with major global representations like Texture, verifying the compatibility and potential of local prompt optimization. It is notable that the extension process is seamless and can be integrated with any global optimization prompts, strongly showing the extensibility of our approach. We also provide density map in Fig. 3 to illustrate the superiority of our method.
ID accuracy
In addition to OOD performance, it is also meaningful to evaluate ID accuracy as it is expected to ensure classification accuracy and separate outliers at the same time. However, weighing too much about separating outliers may narrow the inter-class distance, thus harming accuracy. Consequently, we evaluate ID accuracy for ImageNet-1k and the results shown in Tab. 3 conclude that our method significantly surpasses previous methods (1.02% and 2.38%, respectively). It demonstrates that the proposed regional enhancement is also beneficial to improve OOD separation while increasing inter-class distance, which is of great significance for real application.
Table 4: Experiments on different datasets. ImageNet-100 is used as ID dataset. 4-shot results are reported for few-shot tuning.
Method iNaturalist SUN Places Texture Average
[…]
MCM [16] 18.13 96.77 36.45 94.54 34.52 94.36 41.22 92.25 32.58 94.48
LoCoOp [18] 44.94 93.31 30.70 94.70 34.31 93.93 55.92 90.01 41.47 92.99
Ours 6.99 98.05 24.34 96.03 26.04 95.49 36.32 93.67 23.42 95.81
Table 5: Effectiveness of each component in loss function.
Loss iNaturalist SUN Places Texture Average
[…]
Baseline ($\mathcal{L}_{\mathrm{pos}}$) 11.64 97.59 26.81 94.59 34.50 91.59 50.46 87.84 30.85 92.90
$+\mathcal{L}_{\mathrm{neg}}$ 8.62 98.04 23.82 95.02 32.08 92.12 49.84 88.49 28.59 93.42
$+\mathcal{L}_{\mathrm{neg}}+\mathcal{L}_{\mathrm{reg}}$ 9.39 97.97 24.20 95.10 32.13 92.21 49.37 88.66 28.77 93.48
Near OOD detection tasks
Near OOD task typically evaluates semantically similar categories and is relatively more challenging. Following existing work [16], we conduct experiments on ImageNet-10 vs. ImageNet-20 to verify the effectiveness of the proposed approach. Results in Tab. 3 indicate that our method exceeds previous methods by a large margin. It can be observed that previous few-shot method fails in near OOD tasks. We attribute it to the refinement of local prompts by local outlier knowledge, which greatly assists in distinguishing semantically similar categories. By contrast, relying solely on global prompts may confuse the distribution. Therefore, it underlines that our local outlier enhancement approach is especially effective in semantically similar near OOD tasks.
ImageNet-100 as ID dataset
We also conduct experiments using ImageNet-100 as ID dataset. We compare with other zero-shot and few-shot methods and results are reported in Tab. 4. The substantial performance improvements (average 9.16% reduction on FPR95 and 1.33% promotion on AUROC) strongly demonstrate the effectiveness and transferability of our approach.
5.3 Ablation Study
We conduct various and comprehensive ablation studies to investigate the effectiveness of each component. We train a 4-shot model to showcase the function of components and hyper-parameters unless otherwise stated.
Table 6: Influence of different negative augmentation strategies.
iNaturalist SUN Places Texture Average
[…]
Baseline 11.18 97.68 23.30 95.25 30.97 92.44 50.44 87.89 28.97 93.31
+cutout 17.47 96.64 26.17 94.73 34.38 91.84 54.84 86.88 33.21 92.52
+random crop 9.39 97.97 24.20 95.10 32.13 92.21 49.37 88.66 28.77 93.48
Table 7: Influence of different OOD score strategies. LNP stands for local negative prompts.
Method iNaturalist SUN Places Texture Average
[…]
MCM [16] 27.70 94.66 31.30 93.68 39.50 90.86 38.61 91.79 34.27 92.75
GL-MCM [17] 9.69 97.80 26.27 94.22 34.78 91.33 36.63 91.65 26.84 93.75
Ours(w/o LNP) 7.71 98.18 23.78 94.98 31.85 92.28 35.76 91.79 24.77 94.31
Ours(w LNP) 8.63 98.07 23.23 95.12 31.74 92.42 34.50 92.29 24.52 94.48
Effectiveness of loss components
We conduct experiments to analyse the effectiveness of each loss component and results in Tab. 5 imply that local prompts optimization (baseline) achieves comparable results with global prompts optimization methods, strongly demonstrating the effectiveness of our proposed regional enhancement strategy. Furthermore, taking hard negative samples together with local negative prompts $\bm{t}_{\mathrm{neg}}$ to optimize OOD-considered representation is beneficial to the distinction of outliers. Outcome without regularization loss is better in FPR95 but slightly inferior in AUROC. Considering overall performance and the fact that the optimization of $\bm{t}_{\mathrm{neg}}$ relies heavily on the initialization without regularization, we reserve regularization loss for stability and convergence.
Influence of negative augmentation strategy
We conduct experiments to analyse the influence of negative augmentation strategy. In Tab. 6, baseline does not employ augmentation. Cutout and random crop are used as augmentation, respectively. It is concluded that improper augmentation is unfavorable to detection ability. By contrast, random crop is helpful to enhance regional-related representation and distinguish outliers, which confirms our analysis above.
OOD score strategy
We compare our proposed OOD score with existing scores [16, 17] on 16-shots. We also remove local negative prompts in our OOD score to analyse their function. Results in Tab. 7 imply that our regional OOD score attains better performance than existing strategy, displaying the strength of regional enhancement. Moreover, the performance is further promoted when negative prompts are introduced. We attribute it to possible outlier leverage through negative prompts.
Table 8: Influence of different number of local negative prompts.
$N_{neg}$ iNaturalist SUN Places Texture Average
[…]
100 9.02 98.00 24.71 94.82 32.77 91.87 48.81 88.50 28.82 93.30
200 9.65 97.90 22.58 95.26 32.03 92.18 49.85 88.00 28.52 93.33
300 9.39 97.97 24.20 95.10 32.13 92.21 49.37 88.66 28.77 93.48
Number of local negative prompts
The number of local negative prompts $N_{\mathrm{neg}}$ stands for local outlier knowledge the model learns from negative augmentation. Results in Tab. 8 indicate that the performance is positively related to $N_{\mathrm{neg}}$. However, larger $N_{\mathrm{neg}}$ requires more computational resources. Considering both computational cost and performance, the number 300 is selected.
5.4 Visualization of local prompts
We give a detailed explanation of local prompts and visualize the regions that attract attention from local prompts and local negative prompts. Visualization is shown in Fig. 4 in the order of original picture, ID-related, and several typical OOD-related areas. It implies that local prompts concentrate on ID regions and local negative prompts are helpful to extract ID-irrelevant, i.e., outlier areas corresponding to certain semantics. For instance, local negative prompts successfully focus on sky, sea, and beach, respectively in pirate ship, which is beneficial to enhance OOD detection ability through local and unknown knowledge leverage.
Figure 4: Visualization of ID and OOD related regions. Local prompts and local negative prompts successfully focus on ID-related and OOD-related regions, respectively, which helps OOD detection.
6 Conclusion
In this paper, we investigate the relationship between global and local prompts for few-shot OOD detection. Concentrating on enhancing outlier knowledge for better local textual representation utilization, we decompose global and local features using respective prompts. We use straightforward negative augmentation to leverage local outlier knowledge and propose local prompts to enhance regional information by means of regional regularization. We also propose regional-related evaluation metric to enrich the detector with local outlier information. Moreover, our approach is extensible to well-trained global prompts to get better performance. We conduct various experiments and superior results underline the effectiveness of our method. We hope our method can provide a new perspective on studying local prompts for OOD detection tasks based on vision-language models.